Sol Summary – April 15th

Crew 210 Sol Summary Report 15-APR-2019

Sol: 3

Summary Title: Squashing plants and supply runs

Author’s name: Paul Sokoloff

Mission Status: Nominal

Sol Activity Summary: This morning, after the requisite coffee, Crew 210 pressed the plants collected during yesterdays excursion to Salt Creek and the Burpee Quarry. By squishing the plant specimens between cardboard and drying them out, the team is left with a flat plant that, of stored correctly, will last for hundreds of years. These specimens, accompanied by a label indicating the location and date of collection, will serve as the proof that these plants were found growing at a specific place and time, and will make up the backbone of our first planned publication from this program.

Our team devoted the rest of this sol to gathering supplies essential to hab and rover systems. Shannon made a water run to Hanksville, while the rest of 210 went to Bicknel for ATV batteries and supplies to build a quadrat for our ecological work. On our return to MDRS our team was able to repair one more ATV, three of these are now operational.

While completing both our scientific and technical work for the day we were also able to continue our social outreach efforts, including posting video of our plant pressing adventures to Twitter and Facebook, preparing for an “Ask a Martian Anything” on Reddit tomorrow night, and telling our “Martian” tales on social media.

All in all, another beautiful day in that special place that straddles the line between our unique planet and the one we’d next like to visit.

Look Ahead Plan: Now equipped with a square-meter quadrat for ecological sampling, we will plan on spending tomorrow (sol 4), doing transects and quadrats within the MDRS operational area. At these sites we will document the number of species, percent cover, and functional groups of the plant and lichen communities, and record abiotic variables such as soil moisture content, pH, and coarseness. New plant and lichen species found on EVA will also be collected.

A Bureau of Land Management representative is scheduled to visit MDRS tomorrow morning, so we will meet with them before starting our sampling EVA.

Anomalies in work: No.

Weather: Cloudy in the am, clearing by mid-day.

Crew Physical Status: Good.

EVA: no EVAs at MDRS, but Shannon made a water and supply run to Hanksville, and the remainder of the crew made a supply run to Bicknell.

Reports to be filed:

Sol Summary

Operations Report

GreenHab Report

Photo Report

Support Requested: n/a

Copyright © The Mars Society. All rights reserved. | Main Site